# Predictive maintenance solution for industrial systems - an unsupervised approach based on log periodic power law

### v2

@@ V Data description
Monitoring data from a reciprocating compressor, describing the PV diagram of one of the compression chambers, was used to demonstrate the operation of the method. Based on these, the values of the angle of opening of the suction valve (OSV) expressed by the angle of rotation of the crankshaft were determined.
This value, for a given cycle described by the PV diagram, is very sensitive to changes in the amount of gas in the chamber, for example, due to its leakage through a broken valve or piston rod seal system. OSV changes are directly dictated by the thermodynamics of the compression process in the compressor chamber. Monitoring the changes in OSV provides a basis for compressor diagnostics and allows to determine compressor efficiency, valve operation, or the condition of the piston seals or piston rod sealing elements Reciprocating_compressors .
@@ VI Prediction of failures: methodology and results
To test the effectiveness of our algorithm, a backtest of the detection method was conducted on historical data starting from the initial date of 2019-10-09 ($t_{start}$) to 2022-01-01, calculating initial breakdown (IB) points for this time period. The range of the variable length of the time series $l_{max}$ was assumed to be $30<l_{max}<101$ in time units of days. Given the minimum number of observations (101 days) needed to perform calculations of the IB points, the backtest is started for time $t=t_{start}+101$ (in daily units). Then, moving forward in time to the future, the best-fit LPPL function (26) is calculated for each subsequent time $t$ by determining the goodness of fit of the LPPL function using the mean squared error (msemse).
Intuitively, one can expect that the accuracy of determining the critical points using theorem (1) will strongly depend on the error of fitting the LPPL curve (26) to the data. The smaller the fitting error, the greater the confidence that a given point of the input time series determined as an IB point according to the theorem (1) is really the IB critical point. In addition, it is expected that in the vicinity of the true IB breakpoint (before and after it), the method should find some good fits of the LPPL function with a small error, but this is not a necessary criterion for the existence of an IB point for days as time units.
The figure (1) shows 2 examples of the fit function (26) to data along with calculated criteria for trends determined from maxima and minima of the fitting function satisfying the criteria of Theorem (1).
@@
Figure 1: Examples of LPPL function (26) fits for cases with positive trends (a) and with negative trends (b) for selected current times (dates in proleptic Gregorian ordinal units). Some fit parameters calculated for both trend categories: (a) positive trend: current date: $t_{c}=2020-08-28$, length of the time series: $l_{max}=79$, mean squared error: $mse=2.6\cdot 10^{-5}$; (b) negative trend: current date: $t_{c}=2020-09-23$, length of the time series: $l_{max}=55$, mean squared error: $mse=5.7\cdot 10^{-5}$.
The figure (2) shows the application of this procedure for the diagnosis of the critical points with additional information about the dates of failure repairs (see the description of the figure (2)).
Figure 2: Calculated IB points for the analyzed input time series (marked in the figure by a solid line with measurement points). The vertical lines of different colors indicate the diagnosed by the algorithm IB points. For each group of the IB points, the mean value of the matching error (msemse) of the LPPL function (26) is annotated. The black vertical lines indicate the dates of compressor repairs, which typically took place a few to several weeks after the algorithm detected the fault (IB points). The correlation with the diagnosed breakpoints is clearly visible, except for the prediction determined by the largest error (mse=0.000249) for 2020-12-14.
The figure (2) confirms our initial hypothesis very well. It shows:
@@
    groups of points with similar fitting errors at least 1414 days before the time of failure identification (repair is usually performed with additional delay due to the compressor operating conditions).
@@
    Critical event: ($mse<6\cdot 10^{-5}$) severe failure expected, checking the compressor required and preparation for repair,
@@
As shown, sometimes the algorithm detects a larger number of IB points located very close to each other. Such cases were simplified by choosing a single representation (the first IB point of the group) for each group of signals. This was done by assuming that a signal belongs to a group if its distance from the preceding signal is smaller than or equal to 33 days.
@@ VI.1 Determining the time window of predicted failures
The predicted time period of failure occurrence is defined as the interval $\left<n+\frac{l_{max}}{2},n+90\right$>, where $n$ and $l_{max}$ are the indices of the actual time point $x_{n}$ and the parameter defining the length of the input time series $t_{inp}$ used to find the best fit of the LPPL function (26) to a given value of $t_{inp}\left(x_{n},y_{n}\right)$, respectively.
The definition (2) is based on knowledge of the dynamics of the device for which the algorithm parameters have been defined. For other devices, all parameters should be selected based on the dynamics of their behavior. Duration of the time window with predicted failure is assumed to be valid for a certain period of time and is up to 90 days.
In the case of compressors, due to the different criticality of failures, some of them may be accepted for a longer period of time (even several months in the case of valve failures) to wait for a convenient moment of repair.
@@
  • the classification of alerts specified in the definition (1),
@@
  • the definition (2) specifying the expected time window in which the failure will occur,
the raw results shown in figure (2) can be redrawn to a new form, as shown in figure (3). The correlation between predicted failure times and actual repair times, and the time periods when experts detected abnormal compressor behavior is very good for the Critical event and Monitoring event categories. Predictions in the Irrelevant event category were not confirmed by any repair and diagnosis records.
Figure 3: Redrawn prediction of failures by the algorithm in comparison with reparation times and problems detected by experts. The figure shows representations of groups of discovered IB points (colored and continuous single vertical lines) and predicted failure time periods corresponding with them according to definition (2) (corresponding colored areas bounded by vertical dashed lines). The categorization of the criticality of the forecasted problems (definition (1)) is represented by the splitting into 3 separate figures, each for a separate criticality category. The Monitoring event class remains without any marked vertical lines. The algorithm found no prediction for this category. The reparation (maintenance) dates are indicated in the figures by black vertical lines. Time periods when abnormal compressor behavior requiring monitoring and was not qualified for repair by experts are indicated by the gray colour without additional delimiting lines. The intersection of the areas determined by the algorithm as the period of failure occurrence with periods of abnormal behavior of the compressor can be seen for periods starting from 2021-07-15 in the Critical event category. The raw data of the input time series $t_{inp}$ are indicated by the solid blue line with the measurement points ($o$).
@@ VI.2 Root cause of predicted failures
Thus, based on the theorem (1) and a physical interpretation based on the behavior of the time series, the algorithm is able to predict not only the time window of failure, but also the group of parts that may fail. This provides an opportunity to verify the prediction not only on the basis of event times, but also on the basis of identifying the parts that can fail.
To better illustrate the additional information regarding the location of the future failure, the figure (4) shows the same data as in the Figure (3) with additional information about the parts that actually failed, the parts in which experts have observed problems and prognosis of failures predicted by the algorithm. Details are given in the description of the figure (4).
For the entire time period analyzed, 2 cases that deviate from the diagnosis are visible
@@
    the perturbation identified by experts, started in 202111302021-11-30 and identified as indiation of discharge valve leakage was not predicted by the algorithm at all.
@@
Figure 4: The same graphs as in the figure (3) with added annotations describing the failures predicted by the algorithm - colored texts (different from gray and black), reasons for compressor reparations - black text and diagnosed abnormal compressor behavior requiring monitoring - gray text. For better visualization, areas of the diagnosed abnormal behaviour of the compressor that require monitoring are displayed in the "Monitored Event" category.
@@
Table 1: Detailed comparison of model predictions with maintenance logs entries and expert diagnoses. The "Predictions" section contains a description of the alert: "Alert date" - date of alert occurrence, "mse" - goodness of the fit, "Predicted failure time window" - the predicted time window in which the failure may occur, and "Predicted failure" - the predicted failure diagnosis. The "Maintenance logs" section contains columns: "Maintenance date" - repair date and "Diagnosis" - the reason for the failure. The "Expert diagnoses" columns contain information in turn about: the beginning date of the anomaly notice - "Start date", the end date of the anomaly - "End date" and the symptom of the anomaly - "Diagnosis". The "Label" column contains the identification of the predicted failure. The following convention is used: True Positive (TP) event, with correctly predicted compressor failures or misbehavior, False Positives (FP), an event predicted by the algorithm that turned out to be false, and False Negatives (FN), i.e. compressor failures or instabilities that were not predicted.
@@
VII Discussion
The comparison of the information resulting from the failure predictions and comparing it with the knowledge from maintenance logs and expert detection of periods of anomalous compressor operation, including the prediction classification, is provided in Table 1.
Major challenges in the field of industrial application of failure prediction, especially in the unsupervised version, is a number of issues relating to proper identification of failures and behaviors of monitored devices. These difficulties mainly stem from:
  • the large variety in the types of failures,
  • the small number of failures compared to the amount of data,
  • the large variety of behaviors leading to the same type of failure,
The proposed method introduces a new type of procedure, which is based on the search for common functional behavior (26). From the point of view of the numerical values of the analyzed data, the course of the fitted function can be very different for different events - different patterns of the function for different values of fitted parameters. Even when new data appears with values that did not exist in the past, it is possible to determine the critical point (IB) for a potential event, as the model fits functions to the data.
The IB points determined by this method, upon which the time window of failure occurrence is predicted in the next step, are the trend change points in the data. From this perspective, the described solution can be reduced only to the task of determining the trend change points.
To calculate the key performance indicators (KPIs) of the presented algorithm, the generally known indicators can be used: $Precision=TP/\left(TP+FP\right)$ and $Recall=TP/\left(TP+FN\right)$, where TPTP - defines True Positive events, i.e. correctly predicted failures or abnormal behavior in the operation of the compressor, FPFP - False Positives, i.e. events predicted by the algorithm that turned out to be false and FNFN - False Negatives, i.e. failures and instabilities of the compressor that were not predicted.
In the analysis of results (section VI), the limits of acceptance of errors of fitting the LPPL function (26) determining the criticality of the predicted events (definition 1) were defined on the basis of the results. Therefore, in order to calculate the PrecisionPrecision and RacallRacall indicators, all results are taken into account without distinguishing them due to the defined criticality.
Comparing the predicted failure periods, taking into account the dates and predicted types of failures with the dates and descriptions of Maintenance logs or recorded faults (Figs. 3 and 4 and the table 1), the calculated values of TPTP, FPFP and FNFN are as follows: $TP=4$, $FP=2$, $FN=1$ . Hence $Precision=0.67$, $Recall=0.8$. Given that, this result takes into account the dating of the predicted failure period along with the prediction of the cause of failure, such an outcome is considered very good.
Advantages of the method:
Disadvantages of the method:
  • data: the method is applicable to data that describe a physical process that degrades/changes due to perturbations introduced by interacting elements. This is because the method searches for behavior characteristic of phenomena in which phase transitions can be observed. Hence, not all data are appropriate for the described method.
  • Matching the function to the data is based on the proper determination of boundaries of the parameters to be matched (26). This requires individualized adjustment of the ranges of change of these parameters to the monitored device.
  • It is required to select the time step of the input time series in such a way that it is consistent with the dynamic characteristics of the monitored device.
@@ VIII Conclusions
It was shown that failures in the analyzed industrial system can be treated as critical behavior in complex systems. Then the symptoms of future failure appear in the form of Log Periodic Power Law structures for phase transitions in the analyzed time series. Based on the most generalized scheme for describing the behavior of an analyzed system in the vicinity of phase transitions of the 2nd kind based on the Log Periodic Power Law, a new way of predicting failures in compressor systems is proposed.
@@
Taking into account the specificity of problem detection in industrial systems (the demand to reduce the number of false alarms and to minimize the number of unpredicted events), it has been demonstrated that it is possible to experimentally determine such an error threshold of fitting the LPPL function to the data that all serious failures can be predicted if the fitting error is smaller than the threshold. In addition, it is also possible to define such thresholds for the LPPL curve-fit error to the data, for which the area of occurrence of less critical failures that do not require rapid intervention can be defined.

### v3

@@ V Data description
Monitoring data from a reciprocating compressor, describing the PV diagram of one of the compression chambers, was used to demonstrate the operation of the method. Based on these, the values of the angle of opening of the suction valve (OSV) expressed by the angle of rotation of the crankshaft were determined. The data is available at data_private .
This value, for a given cycle described by the PV diagram, is very sensitive to changes in the amount of gas in the chamber, for example, due to its leakage through a broken valve or piston rod seal system. OSV changes are directly dictated by the thermodynamics of the compression process in the compressor chamber. Monitoring the changes in OSV provides a base for compressor diagnostics and allows to determine compressor efficiency, valve operation, or the condition of the piston seals or piston rod sealing elements Reciprocating_compressors .
@@ VI Prediction of failures: methodology and results
To test the effectiveness of our algorithm, a backtest of the detection method was conducted on the OSV data, calculating initial breakdown (IB) points. The range of the variable length of the time series $l_{max}$ was assumed to be $30<l_{max}<101$ in time units of days. Given the minimum number of observations (101 days) needed to perform calculations of the IB points, the backtest is started for time $t=t_{start}+101$ (in daily units). Then, moving forward in time to the future, the best-fit LPPL function (26) is calculated for each subsequent time $t$ by determining the goodness of fit of the LPPL function using the mean squared error (msemse).
Intuitively, one can expect that the accuracy of determining the critical points using Theorem (1) will strongly depend on the error of fitting the LPPL curve (26) to the data. The smaller the fitting error, the greater the confidence that a given point of the input time series is indeed an IB point according to Theorem (1). In addition, it is expected that in the vicinity of the true IB breakpoint (before and after it), the method should find some good fits of the LPPL function with a small error, but this is not a necessary criterion for the existence of an IB point for days as time units.
Figure (1) shows 2 examples of the fit function (26) to data along with calculated criteria for trends determined from maxima and minima of the fitting function satisfying the criteria of Theorem (1).
@@
Figure 1: Examples of LPPL function (26) fits for cases with positive trends (a) and with negative trends (b) for selected current times (dates in proleptic Gregorian ordinal units). Fit parameters calculated for both trend categories: (a) positive trend: current date: $t_{c}=2020-08-28$, length of the time series: $l_{max}=79$, mean squared error: $mse=2.6\cdot 10^{-5}$; (b) negative trend: current date: $t_{c}=2020-09-23$, length of the time series: $l_{max}=55$, mean squared error: $mse=5.7\cdot 10^{-5}$.
Figure (2) shows the application of this procedure to the diagnosis of the critical points with additional information about the dates of failure repairs (see the description of Figure (2)).
Figure 2: Calculated IB points for the analyzed input time series (marked in figure by a solid line with measurement points). The vertical lines of different colors indicate the diagnosed by the algorithm IB points. For each group of the IB points, the mean value of the matching error (msemse) of the LPPL function (26) is annotated. The black vertical lines indicate the dates of compressor repairs, which typically took place a few to several weeks after the algorithm detected the fault (IB points). The correlation with the diagnosed breakpoints is clearly visible, except for the prediction determined with the largest error (mse=0.000249) for 2020-12-14.
Figure (2) confirms our initial hypothesis very well. It shows:
@@
    groups of points with similar fitting errors at least $14$ days before the time of failure identification (repair is usually performed with additional delay due to the compressor operating conditions),
@@
    Critical event: ($mse<6\cdot 10^{-5}$) severe failure expected, checking of the compressor required and preparation for repair,
@@
As shown, sometimes the algorithm detects a larger number of IB points located very close to each other. Such cases were simplified by choosing a single representation (the first IB point of the group) for each group of signals. This was done by assuming that the signal belongs to a group if its distance from the preceding signal is smaller than or equal to $3$ days.
@@ VI.1 Determining the time window of predicted failures
The predicted time period of a failure occurrence is defined as an interval $\left<n+\frac{l_{max}}{2},n+90\right$>, where $n$ and $l_{max}$ are the indices of the actual time point $x_{n}$ and the parameter defining the length of the input time series $t_{inp}$ used to find the best fit of the LPPL function (26) to a given value of $t_{inp}\left(x_{n},y_{n}\right)$, respectively.
Definition (2) is based on a knowledge of the dynamics of the device for which the algorithm parameters have been defined. For other devices, all parameters should be selected based on the dynamics of their behavior. Duration of the time window with predicted failure is assumed to be valid for a certain period of time and is up to 90 days.
In a case of compressors, due to the different criticality of failures, some of them may be accepted for a longer period of time (even several months in the case of valve failures) while waiting for a convenient moment for repair.
@@
  • the classification of alerts specified in Definition (1),
@@
  • Definition (2) specifying the expected time window in which the failure will occur,
the raw results shown in Figure (2) can be redrawn to a new form, as shown in Figure (3). The correlation between the predicted failure times and the actual repair times, or the time periods when experts detected abnormal compressor behavior is very good for the Critical event and Monitoring event categories. Predictions in the Irrelevant event category were not confirmed by any repair and diagnosis records.
Figure 3: Redrawn prediction of failures by the algorithm in comparison with reparation times and problems detected by experts. Figure shows representations of groups of discovered IB points (colored and continuous single vertical lines) and predicted failure time periods corresponding to them according to Definition (2) (corresponding colored areas bounded by vertical dashed lines). The categorization of the criticality of the forecasted problems (Definition (1)) is represented by the splitting into 3 separate figures, each for a separate criticality category. The Monitoring event class remains without any marked vertical lines. The algorithm found no prediction for this category. The reparation (maintenance) dates are indicated in figures by black vertical lines. Time periods of abnormal compressor behavior requiring monitoring and not qualified for repair by experts are indicated by the gray colour without additional delimiting lines. The intersection of the areas determined by the algorithm as a period of failure occurrence with periods of abnormal behavior of the compressor can be seen for periods starting from 2021-07-15 in the Critical event category. The raw data of the input time series $t_{inp}$ are indicated by the solid blue line with the measurement points ($o$).
@@ VI.2 Root cause of predicted failures
Thus, based on Theorem (1) and a physical interpretation of the behavior of the time series, the algorithm is able to predict not only the time window of failure, but also the group of parts that may fail. This provides an opportunity to verify the prediction not only on the basis of event times, but also on the basis of identifying the parts that can fail.
To better illustrate the additional information regarding the location of the future failure, Figure (4) shows the same data as Figure (3) with additional information about the parts that actually failed, the parts in which experts have observed problems and prognosis of failures predicted by the algorithm. Details are given in the description of Figure (4).
For the entire time period analyzed, 2 cases deviating from this diagnosis are visible
@@
    the perturbation identified by experts, started on 202111302021-11-30 and identified as indiation of discharge valve leakage was not predicted by the algorithm at all.
@@
Figure 4: The same graphs as in Figure (3) with added annotations describing the failures predicted by the algorithm - colored texts (different from gray and black), reasons for compressor reparations - black text and diagnosed abnormal compressor behavior requiring monitoring - gray text. For better visualization, areas of the diagnosed abnormal behaviour of the compressor that require monitoring are displayed in the "Monitored Event" category.
VI.3 Comparison with statistical method
The described method generates IB points before the appearance of anomalies associated with a future failure (section IV). This is why the LPPL-based solution has an advantage over solutions based on anomaly detection as a signature of a forthcoming failure.
It cannot be excluded that the IB points detected by the described method, may be the same as the statistically defined change points in the input trend of the time series. For comparison of the results determined by the discussed solution with the trend change points determined statistically, the online changepoint detection method from the work of changepoint2 was used. This solution is available in the form of the Python library changepoint3 (changepoint_online).
I ran a model based on detection of two states of the analyzed data: one focusing on the detection of a decreasing trend, and the other focusing on the changes in the increasing trend. This way, it is possible to track when the analyzed variable increases and when it decreases. This corresponds to the ability to detect the type of future failure, as described in subsection VI.2.
In each iteration of the historical test, two changepoint detectors are run, one focusing on decreasing changes (left detector) and the other on increasing changes (right detector). In the source of the changepoint_online changepoint3 , this corresponds to the terminology: left and right detectors, respectively. The two detectors work independently of each other. When either model exceeds a threshold value, it is reinitialized.
A comparison of the results of the two methods at the level of determining IB points and trend change points is shown in Figure 5.
Figure 5: Comparison of the results of the presented method (LPPL method - top figure) with the statistical method (changepoint_online - middle and bottom figures) at the level of determining IB points and trend change points. The analyzed time series $t_{inp}$ is shown by the solid blue line with the measurement points ($o$). The black vertical lines indicate the compressor repair dates. Time periods when abnormal compressor behavior requiring monitoring was not qualified for repair by experts are indicated by the gray colour without additional delimiting lines. Top figure: for the LPPL method, the vertical lines in different colors indicate the IB points diagnosed by the algorithm for the LPPL method (as in Figure 2). Middle and bottom figure: the found points of upward (middle figure) and downward trend change (bottom figure). Calculations of trend change points are shown for the threshold value of the 75th percentile.
As shown in Figure 5, the set of IB points determined by the LPPL method (14 alerts) is much smaller than the set of trend change points found by he online statistical method (changepoint_online) (44 left and 33 right detections) . This is consistent with intuition because not every trend change point is a signature of change due to device degradation. The selection of all trend change points will create a lot of false signals. Therefore, it is necessary to correctly identify the threshold value to be used for selection of the best trend change points, which may suggest future failure. The choice of the threshold value should demonstrate the relation between changes in the behavior of the analyzed variable with degradation processes in the monitored device. Such a correlation requires a separate analysis and knowledge of the labeled data. The applied statistical method is unable to predict future failures based solely on determining trend change points in a fully unsupervised manner.
VII Discussion
The comparison of information coming from the failure predictions and comparing it with the knowledge from maintenance logs and expert detection of periods of anomalous compressor operation, including the prediction classification, is provided in Table 1.
Major challenges in the field of industrial application of failure prediction, especially in the unsupervised version, is the number of issues corresponding to proper identification of failures and behaviors of monitored devices. These difficulties mainly stem from:
  • the large variety in the types of failures,
  • the small number of failures compared to the amount of data,
  • the large variety of behaviors leading to the same type of failure,
The proposed method introduces a new type of procedure, which is based on the search for common functional behavior (26). From the point of view of the numerical values of the analyzed data, the course of the fitted function can be very different for different events - different patterns of the function for different values of fitted parameters. Even when new data appears with values that did not exist in the past, it is possible to determine the initial breakdown (IB) for a potential event, as the model fits functions to the data.
The IB points determined by this method, upon which the time window of failure occurrence is predicted in the next step, are the trend change points in the data.
To calculate the key performance indicators (KPIs) of the presented algorithm, the generally known indicators can be used: $Precision=TP/\left(TP+FP\right)$ and $Recall=TP/\left(TP+FN\right)$, where $TP$ - defines True Positive events, i.e. correctly predicted failures or abnormal behavior in the operation of the compressor, $FP$ - False Positives, i.e. events predicted by the algorithm that turned out to be false and $FN$ - False Negatives, i.e. failures and instabilities of the compressor that were not predicted.
In the analysis of results (section VI), the limits of acceptance of errors of fitting the LPPL function (26) determining the criticality of the predicted events (Definition 1) were defined on the basis of the results. Therefore, in order to calculate the $Precision$ and $Racall$ indicators, all results are taken into account without distinguishing them due to the defined criticality.
Comparing the predicted failure periods, taking into account the dates and predicted types of failures with the dates and descriptions of Maintenance logs or recorded faults (Figures 3 and 4 and the table 1), the calculated values of $TP$, FPFP and $FN$ are as follows: $TP=4$, $FP=2$, $FN=1$ . Hence $Precision=0.67$, $Recall=0.8$. Given that, this result takes into account the dating of the predicted failure period along with the prediction of the cause of failure, such an outcome is considered very good.
Advantages of the method:
Disadvantages of the method:
  • The method is applicable to data that describe a physical process that degrades/changes due to perturbations introduced by interacting elements. This is because the method searches for behavior characteristic of phenomena in which phase transitions can be observed. Hence, not all data are appropriate for the described method.
  • Matching the function to the data is based on the proper determination of boundaries of the parameters to be matched (26). This requires individual adjustment of the ranges of change of these parameters and the unit of time in the data to the device being monitored.
The method presented is designed to predict such failures, which are a consequence of a continuous physical process that can be monitored by measurements. In addition, it is necessary to know the time scales specific to the degradation of the monitored equipment. The time unit used by the algorithm should correspond to the time scale units of the degradation process.
@@ VII Discussion
Table 1: Detailed comparison of model predictions with maintenance logs entries and expert diagnoses. The "Predictions" section contains a description of the alert: "Alert date" - date of alert occurrence, "mse" - goodness of the fit, "Predicted failure time window" - the predicted time window in which the failure may occur, and "Predicted failure" - the predicted failure diagnosis. The "Maintenance logs" section contains columns: "Maintenance date" - repair date and "Diagnosis" - the reason for the failure. The "Expert diagnoses" columns contain information about: the beginning date of the anomaly notice - "Start date", the end date of the anomaly - "End date" and the symptom of the anomaly - "Diagnosis". The "Label" column contains the identification of the predicted failure. The following convention is used: True Positive (TP) event, with correctly predicted compressor failures or misbehavior, False Positives (FP), an event predicted by the algorithm that turned out to be false, and False Negatives (FN), i.e. compressor failures or instabilities that were not predicted.
@@ VIII Conclusions
It was shown that failures in the analyzed industrial system can be treated as critical behavior in complex systems. Then the symptoms of future failure in the analyzed time series, appear in the form of Log Periodic Power Law structures characteristic of the vicinity of phase transitions in physical systems. Based on the most generalized scheme for describing the behavior of an analyzed system in the vicinity of phase transitions of the 2nd kind based on the Log Periodic Power Law, a new way of predicting failures in compressor systems is proposed.
@@
Taking into account the specificity of problem detection in industrial systems (the demand to reduce the number of false alarms and to minimize the number of unpredicted events), it has been demonstrated that it is possible to experimentally determine such an error threshold of fitting the LPPL function to the data that all serious failures can be predicted if the fitting error is smaller than the threshold. In addition, it is also possible to define such thresholds for the LPPL curve-fit error to the data, for which the area of occurrence of less critical failures, that do not require rapid intervention, can be defined.
@@
Acknowledgements
The author thank colleagues from Prognost Systems GmbH for the introduction to the topic of failure detection in compressors and for helpful comments and discussions.